OpenStudy (trwatkins1):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (trwatkins1):

hi :)

OpenStudy (youngstudier):

What do you think?

OpenStudy (trwatkins1):

c?

OpenStudy (youngstudier):

C and ?

OpenStudy (trwatkins1):

d?

OpenStudy (youngstudier):

Yep.

OpenStudy (trwatkins1):

i dont understand closed? that just mean zero?

OpenStudy (youngstudier):

-1 + 1 = 0 - Not in the set -1 - 1 = -2 - not in the set. If the result of the action is in the set given, it is closed. If it isn't, it is not closed. -1/1 = -1 - Is in the set, so therefore it is closed -1*1 = -1 - is in the set, so therefore it is closed.

OpenStudy (trwatkins1):

ok...

OpenStudy (youngstudier):

I know, it's tricky, I learned this last year, but it didn't make sense until this year. :P

OpenStudy (trwatkins1):

yeah when i was in first grade my teacher held me back because they were lieing to the state

OpenStudy (youngstudier):

Ah okay.